OpenStudy (anonymous):

You have $22 in your bank account and deposit $11.50 each week. At the same time your cousin has $218 but is withdrawing $13 each week. When will your accounts have the same balance? (Hint: Write a system for the situation and solve)

OpenStudy (anonymous):

22+11.50x=218-13x 11.50x+13x=218-22 24.5x=196 x=8